# Unifying Autoregressive and Diffusion-Based Sequence Generation

### v1

@@ G.1 Undonditonal Sample
Figure 13: Unconditional samples generated by $\gamma$-Hybrid [$\gamma{=0.01},\boldsymbol{\tau}_{\text{Slide}}^{\omega{=\nicefrac{{d}}{{4}}}},\textsc{aligned}$] trained on OWT dataset.
{mdframed}
[ linewidth=1mm, linecolor=black, backgroundcolor=mycustomcolor, roundcorner=5pt ] White House chief of staff Reince Priebus traveled to Washington, D.C., on Wednesday to be briefed by Vice President Mike Pence and other top advisers about the nature of the crisis in Ukraine and its potential impact on the United States.
The new official, who was not authorized to speak publicly by the White House, met with senior American Russian policy experts and other world leaders as part of a trip to Moscow to meet with Ukrainian leader Mykola Azarov. The trip comes as tensions peak over President Donald Trump’s handling of the refugee crisis and its potential for conflict with Russia.
Jared Kushner’s efforts to pump up the Dubrovnik deal have raised concerns that it could backfire. Kushner Trump is accused of colluding with the Kremlin, according to a New York Times report .
White House officials said that press secretary Sean Spicer called Kushner a “little stranger” when asked if he was familiar with the concerns swirling around the discussions, according to an account published by NBC News .
“The president-elect agreed to be briefed by my team on Russian meddling in our election process,” Press Secretary Sean Spicer told NBC News in a telephone call, adding that he didn’t have an authority to disclose details of the discussions.
Versions of the meeting have drawn criticism from U.S. politicians, military officials and human rights advocates.
In a separate report, former acting FBI Director Andrew McCabe said he was not aware of any conversations with any Russian officials. He has said he did not know why the advisers were so closely involved in this scheduling conflict.
“I use very preliminary, telephone calls with colleagues both within our own department and overseas, and I have not had conversations with any of them,” McCabe said of the meetings, which included senior officials and a White House official.
The meetings involve joint efforts by the White House and individual individuals with the understanding that such a meeting would not occur, McCabe said, adding that no other member of the senior staff was involved.
“Please, contact the people you know with whom you have the knowledge in this room. Basically, I am asking you to do the work of reaching out to each one of the top officials associated with your Department of State,” McCabe said.
“Since this is tense, I have trying to contact both the President and his staff and conduct the interview,” McCabe said.
Guidance
White House advisers held a series of meetings with top officials of the Department of State and its federal counterparts. However, the White House’s meetings were themselves influenced by events to help the new administration rule out greater involvement in an escalating crisis that has killed more than 200,000 lives since last August.
In speeches delivered to the nation, Trump touted Friday’s meeting on June 20 as another reminder of the importance of new sanctions against Russia. Trump said Russia’s nuclear weapons “are insulting weak countries” and hoped “for a new normal” after a nosedive in the U.S. elections.
“We have a great relationship with Russia, and we continue to look forward to strengthening our relationship with them,” said Trump.
The “warm” Russia policy has deepened the already long-standing divisions over the Obama administration, and Putin’s recent ally Russia, who built a buffer zone along the Black Sea in 2013, ordered an invasion in 2014 of eastern Ukraine.
White House spokesman Sean Spicer said he was “careful” with ties with Russia.¡—endoftext—¿On Tuesday night, LeBron James received news he was going — so often — from the Hall of Fame. Cleveland’s most respected individual, who most known for his love of basketball, was getting a heart-to-heart with one of the most prestigious prospects in pro basketball history, sitting in the stands as the Knicks’ first guard of the year. The official photo for the Knicks’ current superstar, George Conte, was taken by the longtime reality television legend Wade Boggs.
We demand James get a heart. Exactly what he deserves is hard to guess. People to the point a man taken to heart who never weeks ago was the kind of person a champion of truth would want to meet or make close friends with without his inner purpose. After all, people feel very similar if they haven’t seen Wade Boggs’ latest film, Clump, in which he takes matters into his own hands and seizes his wife — the same wife that never had a child up for adoption.
This event has the potential to be a sour revenge for the Kevin Durant coronation that began a week ago – a winner’s assuredly standard season that had little impact on James’ looming future. But that’s not the case here, either; the only one he is saying is that the uninspiring
@@
Figure 14: Unconditional samples generated by $\gamma$-Hybrid [$\gamma{=0.01},\boldsymbol{\tau}_{\text{Block}}^{\omega{=\nicefrac{{d}}{{64}}}},\textsc{shifted}$] trained on OWT dataset.
{mdframed}
[ linewidth=1mm, linecolor=black, backgroundcolor=mycustomcolor, roundcorner=5pt ] be violent might be protected for one of the reasons he challenged.
Tsherberg is talking about former Republican Rep. Joe Burns who, in many ways tried to do better politically, canceled my appearance in 2012. She said that when the Republican president and we were playing with one other’s decisions, it was the Chechen deal.
“Every time I think, there’s nothing in the issue,” she said. “There comes a time when I faced with reality in a very difficult way. It’s not something I want to put the side of the my team past me.”
During O’Reilly Factor in December 2011, when I dealt with most of my viewers, I watched the Obama Britain show again.
“I was watching very, very closely, knowing that there was an opportunity, I knew I didn’t want to be able to be able to help the Americans. I couldn’t help that,” I said.
Cack Obama also said several people without any authority, just due diligence, have commented on his plan to return to Chechnistan and Russia. T. Obama said he’s not against Putin, and referred to himself as Putin-PM, who said he’s taking “America’s values forward.”
“I think that’s a turning point,” Iza said in 2008 just before the start of my comments, referring to Iza’s parent party, Moscow’s Communist Party, in 1997, according to The Associated Press.
Asked whether Iza believes the economy and its security should run well into the Chechen agreement, and the leader ISIS was this month, Iza said that Trump is a far to go, more conservative agenda.
“He can’t be a politician — he’s not a tough guy who doesn’t take a person,” she said. “My people have respect too. My children are not civil andRoll but they do not [acc believe me].”
After speaking at Trump’s Palm Beach, a bus stop in Texas on Sunday, his attorney, Republican Frank Agoodz, said neither he nor Iza accused me of arguing at the “Maxby Restaurant” in Texas.
“Williams, I want them to help, I don’t want them to be able to help me,” Iza was recorded as saying.
“He said, ‘You have an opportunity in doing this and then you have to do this, you can do this; it’s starting, that’s getting the right people to do it, and he said, ‘They’re going and did, yeah,,’” she tells The Associated Press on Friday evening. Sokherberg had not prior known Iza’s release — since his public hearing, he hasn’t has been expecting anything since then.
But Iza said this is a situation that will cause “Great pain” for all his family.
“It’s important to my well-being, to care for others, about peace and just to be strong,” she said. “As long as I don’t, I won’t. I think if I won’t, I will have no apologies.”
Their representatives before they gather on Saturday.
“It is a man with the very strong side of things, that has a long way to go,” she said. “He’s been singled out by the President. He has turned even more extreme off the map and is being led by the Republican party and the economy that are making a debt-laden economy out of his dream.
“The public is a mind-set, and his decision to speak to the public was a mind-set.”
In a CNN video excerpt that includes her closing remarks, the then-U.S. Senate wasift to safety to stand in front of Republican House-time House Speaker and former White House special counsel Paul Pro Romanoff.¡—endoftext—¿New Villa in the midst of not making a cause in Sunday’s clash at Villa, who started toestern to the title last season.
With Villa attacking you no certainly doubt it, if it was the second title they had won in the most recent five years, over the likes of Marc Canado, Robbie, Fernandes and CFOi.
But Villa think an answer to that battle over Canado is the help of the Vy’s and in speeding up the relationship with the Portuguese forward in ways he hasn’t shown in his past season.
West Villa turned their attention to defence in with Canado, who represented the England national team even last season after he became the first ever to score four goals.
He went on the same defensive side though, is now holding up front to rival one of the top midfielders and attack in the second half.
West Villa’s second-half form last season however, has been done again in favour to substitute Scott-Fidane Jr, who later moved to Paris City on Wednesday.
John Gomez and João Aguero have moved to
@@
Figure 15: Unconditional samples generated by $\epsilon$-Hybrid [$\epsilon{=0.01},\boldsymbol{\tau}_{\text{Flat}},\textsc{aligned}$] trained on OWT dataset.
@@ G.2 Conditional Samples
{mdframed}
[ linewidth=1mm, linecolor=black, backgroundcolor=mycustomcolor, roundcorner=5pt ] The cat sat on the mat and looked up, staring at the ceiling as if something interesting were up there. It made no sound, but occasionally flicked its tail back and forth.
“Is he always like this?” I asked.
She smiled and nodded. “Pretty much. Sometimes he stares out the window too.”
The cat glanced at me, seemed to consider something deeply philosophical for a moment, and then resumed staring upward.
“Do you think he sees things we don’t?” I asked, half-joking.
“Probably,” she replied, laughing softly. “Cats always seem to have a foot in another world, don’t they?”
I chuckled, taking a sip from my coffee. “Maybe we should take notes.”
“Maybe we should,” she said, still smiling. “We might learn something.”
The cat yawned lazily, stretched, and settled even more comfortably onto the mat, clearly deciding that whatever secrets the universe held could wait a little longer.
Figure 16: Conditional (conditioned on first 6 tokens) samples generated by $\epsilon$-Hybrid [$\epsilon{=0.01},\boldsymbol{\tau}_{\text{Flat}},\textsc{aligned}$] trained on OWT dataset.

### v2

@@ G.1 Undonditonal Sample
Figure 12: Unconditional samples generated by $\gamma$-Hybrid $\gamma{=0.01},\boldsymbol{\tau}_{\text{Slide}}^{\omega{=\nicefrac{{d}}{{4}}}},\textsc{aligned}$] trained on OWT dataset.
White House chief of staff Reince Priebus traveled to Washington, D.C., on Wednesday to be briefed by Vice President Mike Pence and other top advisers about the nature of the crisis in Ukraine and its potential impact on the United States. The new official, who was not authorized to speak publicly by the White House, met with senior American Russian policy experts and other world leaders as part of a trip to Moscow to meet with Ukrainian leader Mykola Azarov. The trip comes as tensions peak over President Donald Trump’s handling of the refugee crisis and its potential for conflict with Russia. Jared Kushner’s efforts to pump up the Dubrovnik deal have raised concerns that it could backfire. Kushner Trump is accused of colluding with the Kremlin, according to a New York Times report . White House officials said that press secretary Sean Spicer called Kushner a “little stranger” when asked if he was familiar with the concerns swirling around the discussions, according to an account published by NBC News . “The president-elect agreed to be briefed by my team on Russian meddling in our election process,” Press Secretary Sean Spicer told NBC News in a telephone call, adding that he didn’t have an authority to disclose details of the discussions. Versions of the meeting have drawn criticism from U.S. politicians, military officials and human rights advocates. In a separate report, former acting FBI Director Andrew McCabe said he was not aware of any conversations with any Russian officials. He has said he did not know why the advisers were so closely involved in this scheduling conflict. “I use very preliminary, telephone calls with colleagues both within our own department and overseas, and I have not had conversations with any of them,” McCabe said of the meetings, which included senior officials and a White House official. The meetings involve joint efforts by the White House and individual individuals with the understanding that such a meeting would not occur, McCabe said, adding that no other member of the senior staff was involved. “Please, contact the people you know with whom you have the knowledge in this room. Basically, I am asking you to do the work of reaching out to each one of the top officials associated with your Department of State,” McCabe said. “Since this is tense, I have trying to contact both the President and his staff and conduct the interview,” McCabe said. Guidance White House advisers held a series of meetings with top officials of the Department of State and its federal counterparts. However, the White House’s meetings were themselves influenced by events to help the new administration rule out greater involvement in an escalating crisis that has killed more than 200,000 lives since last August. In speeches delivered to the nation, Trump touted Friday’s meeting on June 20 as another reminder of the importance of new sanctions against Russia. Trump said Russia’s nuclear weapons “are insulting weak countries” and hoped “for a new normal” after a nosedive in the U.S. elections. “We have a great relationship with Russia, and we continue to look forward to strengthening our relationship with them,” said Trump. The “warm” Russia policy has deepened the already long-standing divisions over the Obama administration, and Putin’s recent ally Russia, who built a buffer zone along the Black Sea in 2013, ordered an invasion in 2014 of eastern Ukraine. White House spokesman Sean Spicer said he was “careful” with ties with Russia.¡—endoftext—¿On Tuesday night, LeBron James received news he was going — so often — from the Hall of Fame. Cleveland’s most respected individual, who most known for his love of basketball, was getting a heart-to-heart with one of the most prestigious prospects in pro basketball history, sitting in the stands as the Knicks’ first guard of the year. The official photo for the Knicks’ current superstar, George Conte, was taken by the longtime reality television legend Wade Boggs. We demand James get a heart. Exactly what he deserves is hard to guess. People to the point a man taken to heart who never weeks ago was the kind of person a champion of truth would want to meet or make close friends with without his inner purpose. After all, people feel very similar if they haven’t seen Wade Boggs’ latest film, Clump, in which he takes matters into his own hands and seizes his wife — the same wife that never had a child up for adoption. This event has the potential to be a sour revenge for the Kevin Durant coronation that began a week ago – a winner’s assuredly standard season that had little impact on James’ looming future. But that’s not the case here, either; the only one he is saying is that the uninspiring
@@
Figure 13: Unconditional samples generated by $\gamma$-Hybrid $\gamma{=0.01},\boldsymbol{\tau}_{\text{Block}}^{\omega{=\nicefrac{{d}}{{64}}}},\textsc{shifted}$] trained on OWT dataset.
be violent might be protected for one of the reasons he challenged. Tsherberg is talking about former Republican Rep. Joe Burns who, in many ways tried to do better politically, canceled my appearance in 2012. She said that when the Republican president and we were playing with one other’s decisions, it was the Chechen deal. “Every time I think, there’s nothing in the issue,” she said. “There comes a time when I faced with reality in a very difficult way. It’s not something I want to put the side of the my team past me.” During O’Reilly Factor in December 2011, when I dealt with most of my viewers, I watched the Obama Britain show again. “I was watching very, very closely, knowing that there was an opportunity, I knew I didn’t want to be able to be able to help the Americans. I couldn’t help that,” I said. Cack Obama also said several people without any authority, just due diligence, have commented on his plan to return to Chechnistan and Russia. T. Obama said he’s not against Putin, and referred to himself as Putin-PM, who said he’s taking “America’s values forward.” “I think that’s a turning point,” Iza said in 2008 just before the start of my comments, referring to Iza’s parent party, Moscow’s Communist Party, in 1997, according to The Associated Press. Asked whether Iza believes the economy and its security should run well into the Chechen agreement, and the leader ISIS was this month, Iza said that Trump is a far to go, more conservative agenda. “He can’t be a politician — he’s not a tough guy who doesn’t take a person,” she said. “My people have respect too. My children are not civil andRoll but they do not [acc believe me].” After speaking at Trump’s Palm Beach, a bus stop in Texas on Sunday, his attorney, Republican Frank Agoodz, said neither he nor Iza accused me of arguing at the “Maxby Restaurant” in Texas. “Williams, I want them to help, I don’t want them to be able to help me,” Iza was recorded as saying. “He said, ‘You have an opportunity in doing this and then you have to do this, you can do this; it’s starting, that’s getting the right people to do it, and he said, ‘They’re going and did, yeah,,’” she tells The Associated Press on Friday evening. Sokherberg had not prior known Iza’s release — since his public hearing, he hasn’t has been expecting anything since then. But Iza said this is a situation that will cause “Great pain” for all his family. “It’s important to my well-being, to care for others, about peace and just to be strong,” she said. “As long as I don’t, I won’t. I think if I won’t, I will have no apologies.” Their representatives before they gather on Saturday. “It is a man with the very strong side of things, that has a long way to go,” she said. “He’s been singled out by the President. He has turned even more extreme off the map and is being led by the Republican party and the economy that are making a debt-laden economy out of his dream. “The public is a mind-set, and his decision to speak to the public was a mind-set.” In a CNN video excerpt that includes her closing remarks, the then-U.S. Senate wasift to safety to stand in front of Republican House-time House Speaker and former White House special counsel Paul Pro Romanoff.¡—endoftext—¿New Villa in the midst of not making a cause in Sunday’s clash at Villa, who started toestern to the title last season. With Villa attacking you no certainly doubt it, if it was the second title they had won in the most recent five years, over the likes of Marc Canado, Robbie, Fernandes and CFOi. But Villa think an answer to that battle over Canado is the help of the Vy’s and in speeding up the relationship with the Portuguese forward in ways he hasn’t shown in his past season. West Villa turned their attention to defence in with Canado, who represented the England national team even last season after he became the first ever to score four goals. He went on the same defensive side though, is now holding up front to rival one of the top midfielders and attack in the second half. West Villa’s second-half form last season however, has been done again in favour to substitute Scott-Fidane Jr, who later moved to Paris City on Wednesday. John Gomez and João Aguero have moved to
@@
Figure 14: Unconditional samples generated by $\epsilon$-Hybrid $\epsilon{=0.01},\boldsymbol{\tau}_{\text{Flat}},\textsc{aligned}$] trained on OWT dataset.
@@ G.2 Conditional Samples
The cat sat on the mat and looked up, staring at the ceiling as if something interesting were up there. It made no sound, but occasionally flicked its tail back and forth. “Is he always like this?” I asked. She smiled and nodded. “Pretty much. Sometimes he stares out the window too.” The cat glanced at me, seemed to consider something deeply philosophical for a moment, and then resumed staring upward. “Do you think he sees things we don’t?” I asked, half-joking. “Probably,” she replied, laughing softly. “Cats always seem to have a foot in another world, don’t they?” I chuckled, taking a sip from my coffee. “Maybe we should take notes.” “Maybe we should,” she said, still smiling. “We might learn something.” The cat yawned lazily, stretched, and settled even more comfortably onto the mat, clearly deciding that whatever secrets the universe held could wait a little longer.
Figure 15: Conditional (conditioned on first 6 tokens) samples generated by $\epsilon$-Hybrid $\epsilon{=0.01},\boldsymbol{\tau}_{\text{Flat}},\textsc{aligned}$] trained on OWT dataset.